OpenStudy (anonymous):

A cyclist maintains a constant velocity of 4 m/s headed away from point A. At some initial time, the cyclist is 253 m from point A. What will be his displacement from his starting position after 64 s? Answer in units of m What will be his position from point A after that time? Answer in units of m

OpenStudy (anonymous):

For the initial time, the cyclist is 253 m away from point A, calling d0 to this position we have: d0(To point A) = 253 m 64 seconds after (t = 64s), the cyclist is in a different position, lets call it d1, that can be calculated by: \[v = d1/t\] v = velocity d1 = distance t = time Since the cyclist velocity is constant during this period the resulting distance is equal to: \[d1 = v \times t = 4 \times 64 = 256 m\] So, for the final distance, we just need to add this two displacements: \[d(final) = d0 + d1 = 253 + 256 = 509 m\] Answer: 509 m